OpenStudy (anonymous):

while running on his hands, an athlete sprinted 50.0 meters in a record 16.9 seconds. determine the average velocity for this feat

OpenStudy (radar):

Average velocity would be using d=rt (d= distance, r=rate, t= time) 50.0 meter=r*16.9 seconds r=50.0/16.9 mps(meter per sec) solve for r and precision to one tenth.

OpenStudy (anonymous):

which the answer is 2.96 m/s forward right?

OpenStudy (radar):

That is correct, just be sure you round it to the accuracy of the information provided and to the wishes or instructions you've received from your instructor. I personally would use 3.0 or 2.9 because the data provided was only accurate to the tenth.

OpenStudy (anonymous):

Right thanks :)